OpenStudy (anonymous):

write this equation in slope intercept form: 18x+18y=36

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

just solve for y. For the first one subtract 18x from both sides then divide both sides by 18 to get y=2-x. For the second one subtract 13 from both sides and then divide by -4 to get y=(1/2)x + 13/4

OpenStudy (anonymous):

just solve for y. For the first one subtract 18x from both sides then divide both sides by 18 to get y=2-x. sorry just this

OpenStudy (anonymous):

18x+18y=36 -18x -18x 18y=-18x+36 /18 /18 y=-x+2 <------------ TADA